OpenStudy (anonymous):

derivative of x^2 ln(sqrt(2x+1)

OpenStudy (phi):

you should use the product rule d(u v) = u dv + v du and the chain rule

OpenStudy (anonymous):

i have the first half of my equation right but ii'm doubting the second part. this was my final equation= 2x ln(sqrt(2x+1))+ x^2. (4x^2 + 4x + 1)^-1.5

OpenStudy (phi):

yes the first term is ok For the 2nd term , we will use d ln(u) = 1/u du the 2nd term is \[ x^2 \frac{d}{dx} \ln(u) = x^2 \frac{1}{u} \ \frac{d}{dx} u \] where \( u = \sqrt{2x+1} =\left(2x+1\right)^\frac{1}{2} \) we get \[ \frac{x^2}{\left(2x+1\right)^\frac{1}{2}} \frac{d}{dx}\left(2x+1\right)^\frac{1}{2} \] can you finish ?

OpenStudy (anonymous):

i'm seriously lost, the notation kind of confuses me =/

OpenStudy (anonymous):

d/dx = 0.5 (2x+1)^-1.5

OpenStudy (anonymous):

so i'm thinking x^2/(2x+1)^0.5 . 0.5 (2x+1)^-1.5 . 2

OpenStudy (phi):

the derivative of u^n = n u^(n-1) \[ d u^n = n \ u^{n-1} du \] that says: to take the derivative of u to the nth power, multiply u by n, and decrement the exponent by 1. then multiply all of that by the derivative of u (with respect to whatever variable we are using... in this case with respect to x)

OpenStudy (phi):

*** d/dx = 0.5 (2x+1)^-1.5 *** How are you getting an exponent of -1.5 ? you should be doing (n-1) where n is 1/ 2

OpenStudy (anonymous):

ahh good lord, so it should be 0.5 (2x+1)^-1/2 so x^2/(2x+1)^0.5 . 0.5 (2x+1)^-1/2 . 2

OpenStudy (phi):

which can be simplified

OpenStudy (phi):

\[ \frac{x^2}{\sqrt{2x+1} } \cdot \frac{1}{\cancel2} \frac{1}{\sqrt{2x+1} } \cdot \cancel2 \\ =\frac{x^2}{2x+1} \] that is the 2nd term. The full answer is \[ 2x\ \ln\left( \sqrt{2x+1}\right) + \frac{x^2}{2x+1} \]

OpenStudy (anonymous):

sry, just 1 last question, the (x^2/ sqrt(2x+1)) . (1/ sqrt(2x+1)), how does that turn in to just (x^2/ sqrt(2x+1))

OpenStudy (phi):

when you multiply fractions, you multiply top times top and bottom times bottom for the bottom we use \[ \sqrt{x} \cdot \sqrt{x} = \sqrt{x \cdot x} = \sqrt{x^2} = x \]

OpenStudy (phi):

the bottom is not sqr(2x+1) it is just (2x+1)

OpenStudy (anonymous):

oh now i see it, THANK YOU VERY MUCH!!

OpenStudy (phi):

yw